OpenStudy (anonymous):

i need simple equations for finding base, suface area, and volume for pyramids and cubes and cones and cylinders

OpenStudy (anonymous):

For volume in pyramids, you have to use the formula V = 1/3Bh. For cone, V = 1/3pi(r^2)h. For cubes, it's V = s^3. For cylinders, V = pi(r^2)h. For surface area, pyramid: B + 1/2lP. Cone: pi(r^2) + 1/2d(pi)lh. Cube is simply s^2 + 4s^2.